the reaction of finely divided aluminum and iron(III) oxide, Fe2O3, is called the thermite reaction. it produces a tremendous amount of heat, making the welding od railroad track possible. the reaction of 500.0 grams of aluminum and 500.0 grams of iron(iii) oxide produces 166.5 grams iron. calculate the mass of iron that should be released by this reaction.


1
Expert's answer
2021-10-27T06:40:45-0400

Solution:

The molar mass of Al is 26.98 g/mol

The molar mass of Fe is 55.845 g/mol

The molar mass of Fe2O3 is 159.69 g/mol


Calculate moles of each reactant:

(500.0 g Al) × (1 mol Al / 26.98 g Al) = 18.53 mol Al

(500.0 g Fe2O3) × (1 mol Fe2O3 / 159.69 g Fe2O3) = 3.13 mol Fe2O3


Balanced chemical equation:

2Al(s) + Fe2O3(s) → 2Fe(s) + Al2O3(s)


According to the chemical equation above:

2 mol of Al reacts with 1 mol of Fe2O3

Thus, 18.53 mol of Al reacts with:

(18.53 mol Al) × (1 mol Fe2O3 / 2 mol Al) = 9.265 mol Fe2O3

However, initially there is 3.13 mol of Fe2O3 (according to the task).

Therefore, Fe2O3 acts as limiting reagent and Al is excess reagent.


According to stoichiometry:

1 mol of Fe2O3 produces 2 mol of Fe

Thus, 3.13 mol of Fe2O3 produces:

(3.13 mol Fe2O3) × (2 mol Fe / 1 mol Fe2O3) = 6.26 mol Fe


Calculate the mass of Fe:

(6.26 mol Fe) × (55.845 g Fe / 1 mol Fe) = 349.59 g Fe = 349.6 g Fe

The mass of iron (Fe) is 349.6 grams


Answer: 349.6 grams of iron (Fe) should be released by this reaction.